# Searching for GEMS: The Occurrence of Giant Planets orbiting M-dwarfs within 100 pc

### v1

@@ 4.6 Stage 3: Observational and Literature Vetting
Table 2: The 23 objects which survived all auto and manual vetting prior to spectroscopic and literature validation. Of those 23, 1 is a new candidate GEM; 7 are previously confirmed GEMS; and the rest (15) are astrophysical false positives, i.e., not transiting planets. SB1/2/3 refer to single, double, or triple spectroscopic binaries.
@@
243921117 WASP 80 b Confirmed Early Triaud et al. (2013)
46432937 TIC 46432937 b Confirmed Early Hartman et al. (2024)
243641947 TOI-3235 b Confirmed Mid Hobson et al. (2023)
335590096 TOI-4860 b Confirmed Mid Almenara et al. (2023)
419411415 TOI-5205 b Confirmed Mid Kanodia et al. (2023b)
178709444 TOI-762 A b Confirmed Mid Hartman et al. (2024)
67512645 TOI-6894 b Confirmed Late Bryant et al. (2025)
@@
272519426 TOI-5693.01 BD Mid Irwin et al. (2010)
381856446 WASP 145 A b (TIC 381856447 b) Background Planet Mid Hellier et al. (2019)
At this point, we referred back to our sample of known GEMS within 100 pc (see Section 2). We confirmed that we had successfully recovered all seven objects, reinforcing the robustness of our pipeline (see Table 2). Furthermore, we confirmed that we had recovered or reasonably discounted from our analysis all TOIs occurring within our input catalog and search bounds as of August 10, 2025. Only one TOI (5850), was removed prior to Stage 2, due to noisy observational data. This case, alongside other TOIs not explicitly mentioned in the main text of this work, is fully explained in Appendix Table 7.
We then consulted the TESS EB catalog (Prša et al., 2022) as well as previous GEMS and EB surveys to determine the disposition of our remaining 16 candidates. These objects are listed in the bottom section of Table 2. We identified one candidate as an EB (Prša et al., 2022) and one candidate as a BD (Irwin et al., 2010). We identified a further signal, ostensibly from TIC 381856446 / WASP 145 B, as a confirmed hot Jupiter orbiting the target’s much brighter K2 V companion, WASP 145 A (Hellier et al., 2019). The remaining 13 candidates were subjected to spectroscopic validation using HPF, NEID and WINERED (see below for details on the use of these instruments.)
Spectroscopic vetting reduced our new candidate GEMS sample from 13 to 1, and our total remaining sample – including previously confirmed GEMS – from 20 to 8. This reduction of more than 50% highlights the importance of spectroscopic follow-up to eliminate FPs.
Table 3: Selected orbital parameters of all confirmed and candidate GEMS retrieved in our sample. For the new candidate and TOI-5693 (rows 1 and 2), parameters are quoted from our own juliet fits. For all other objects, parameters are quoted from the references in the final column. All stellar radii are taken from Mann et al. (2015). Within each table subsection, objects are sorted by $R_{p}$.
@@
146712644 $1.52^{+7.43\cdot 10^{-7}}_{-7.27\cdot 10^{-7}}$ $1.86^{+0.213}_{-0.146}$ $0.67^{+0.11}_{-0.08}$ $0.303$ This work
178709444 TOI-762 A b $3.47\pm 7.2\cdot 10^{-7}$ $0.744\pm 0.017$ $0.7556^{+0.0057}_{-0.0067}$ $0.428$ Hartman et al. (2024)
335590096 TOI-4860 b $1.52\pm 3.0\cdot 10^{-7}$ $0.77\pm 0.03$ $0.29^{+0.06}_{-0.11}$ $0.355$ Almenara et al. (2023)
243921117 WASP 80 b $3.07^{+8.3\cdot 10^{-7}}_{-7.9\cdot 10^{-7}}$ $0.999^{+0.030}_{-0.031}$ $0.215^{+0.020}_{-0.022}$ $0.636$ Triaud et al. (2013)
243641947 TOI-3235 b $2.59\pm 4.1\cdot 10^{-7}$ $1.02\pm 0.044$ $0.511^{+0.011}_{-0.012}$ $0.371$ Hobson et al. (2023)
419411415 TOI-5205 b $1.63\pm 1.0\cdot 10^{-6}$ $1.03\pm 0.03$ $0.0016\pm 0.0002$ $0.400$ Kanodia et al. (2023b)
46432937 TIC 46432937 b $44\pm 8.7\cdot 10^{-8} 1.1$$19\pm 0.030$ $0.825^{+0.0157}_{-0.0073}$ $0.539$ Hartman et al. (2024)
67512645 TOI-6894 b $3.37\pm 0.000296$ $1.33^{+0.57}_{-0.43}$ $0.927^{+0.551}_{-0.530}$ $0.220$ Bryant et al. (2025)
@@ 4.6.1 HPF – Northern Hemisphere Targets I
For M-dwarfs in the northern hemisphere, we used the Habitable-zone Planet Finder (HPF; Mahadevan et al., 2012, 2014) to obtain observations on our surviving candidates after the aforementioned checks. HPF is a near-infrared (8080127808080-12780 Å), environmentally stabilized (Stefánsson et al., 2016), fiber-fed (Kanodia et al., 2018) spectrograph with laser-frequency comb calibration (Metcalf et al., 2019) on the 10 m Hobby-Eberly Telescope181818 Based on observations obtained with the Hobby-Eberly Telescope (HET), which is a joint project of the University of Texas at Austin, the Pennsylvania State University, Ludwig-Ludwig-Maximilians-Universität München, and Georg-August Universität Gottingen. The HET is named in honor of its principal benefactors, William P. Hobby and Robert E. Eberly. (HET; Ramsey et al., 1998; Hill et al., 2012, 2021) at the McDonald Observatory in West Texas, USA. We used the HxRGproc package (Ninan et al., 2018) to correct for bias, non-linearity, cosmic rays and process the HPF slope images. We then used barycorrpy (Kanodia & Wright, 2018) to perform the barycentric correction on the individual spectra, which is the Python implementation of the algorithms from Wright & Eastman (2014). Each visit consisted of two exposures of 969 seconds each that are subsequently combined by weighted averaging.
We followed the procedure described in Cañas et al. (2023) to cross-correlate the observed spectra with synthetic spectra from the PHOENIX library (Husser et al., 2013) using the algorithms presented in Zucker (2003). We only cross-correlate the HPF spectra in the echelle orders 464-6 ($\lambda\sim 8530-8890$ Å) and 141814-18 ($\lambda\sim 9930-10760$ Å) because of the minimal amount of telluric contamination in those regions. The resultant cross-correlation functions (CCFs) with HPF spectra are shown in Figure 7, and objects with multiple peaks are identified as FPs (hierarchical stellar systems). Using HPF spectra, we identified eight candidates that are spectroscopic binaries and are listed in Table 2. A detailed characterization of these false positives is beyond the scope of this manuscript and is left to future work (e.g., Boone et al., 2025).
Figure 7: CCF plots for the 12 TICs cited in Table 2 as astrophysical false positives in this work. The different colors represent distinct spectroscopic epochs, with CCF shifts indicating Doppler motion inconsistent with a planetary companion. In particular, the presence of multiple peaks or large velocity variations between observations reveal the target as an EB rather than a single-lined star hosting a planet.
@@ 4.6.2 NEID – Northern Hemisphere Targets II
We also used NEID, an environmentally stabilized (Robertson et al., 2019), fiber-fed (Kanodia et al., 2023c) high-resolution optical (3800 – 9300 Å) spectrograph (Schwab et al., 2016) on the 3.5 m WIYN telescope [^19] The WIYN Observatory is a joint facility of the NSF’s National Optical-Infrared Astronomy Research Laboratory, Indiana University, the University of Wisconsin-Madison, Pennsylvania State University, Purdue University and Princeton University. at Kitt Peak, Arizona, USA to spectroscopically validate some of our candidates. Each visit consisted of an 1800 s exposure using high-resolution mode ($R\sim$ 110,000). We used the Level 2 products from the NEID DRPv1.4.0202020https://neid.ipac.caltech.edu/docs/NEID-DRP/, which includes the wavelength-calibrated spectra. The NEID DRP also calculates the CCF for all objects by cross-correlating the observed spectra with binary masks (e.g., Baranne et al., 1996) that are customized to various spectral types (Bender et al., 2022). We used these CCFs to confirm that TIC-95112238 and TIC-60910638 were both spectroscopic double-lined binaries (SB2) with two peaks in the CCF for each star (see Figure 7).
@@ 4.6.3 WINERED – Southern Hemisphere Targets
4.7 New GEMS Candidate – TIC 146712644
Of the 12 candidates for spectroscopic vetting, we observed 11 with the instruments described above. Due to constraints in observability and telescope time, we were unable to obtain spectra for TIC 146712644; we chose to defer its confirmation or elimination to future work, and we retain it as a candidate in this work. We list its system properties alongside those of the seven pre-confirmed GEMS in Table 3.
TIC 146712644 shows evidence of a planetary companion with $R_{p}=1.86^{+0.213}_{-0.146}~R_{J}$ and $P\approx 1.52$ days. TIC 146742645 (hereafter, “the contaminant”) lies within the same TESS pixel as TIC 1467122644 with a separation of 4″ from the target. Although the target lies at $89.97~\pm~0.25$ pc with $T\text{-mag}=14.56$, the contaminant lies at a distance of $1759.63~\pm~114.86$ pc with a $m_{\text{T}}=15.02$. [^21] Distances calculated from Gaia DR3 parallax and parallax uncertainty measurements. The small brightness difference of only $0.46$ magnitudes makes it difficult to determine, using TESS data alone, which object hosts the detected signal.
As discussed in Section 4.4, such cases were forwarded to the juliet vetting stage to perform a density check. The juliet-derived stellar density of the target $\rho_{\text{targ, juliet}}=12.58^{+1.29}_{-1.22}~~\text{g cm}^{-3}$, compared to our accepted value from the relations in Mann et al. (2015) $\rho_{\text{targ, Mann}}=14.19~~\text{g cm}^{-3}$. Meanwhile, the stellar density of the contaminant $\rho_{\text{cont}}=0.77~~\text{g cm}^{-3}$ (from ExoFOP). This disparity suggests that, if the signal comes from a planet and not an EB, it originates from the target instead of the contaminant. Based on this we retain TIC 146712644 as a candidate. However, we acknowledge that future spectroscopic observations are necessary to validate it.
@@ 5.1 Injection and Recovery
  • 0.7 $R_{J}$ $R_{\text{p, inj}}<1.5~R_{J}$, sampled in $R_{p}$ space. [^22] We chose to sample injections in $R_{p}$ as opposed to $R_{p}/R_{\star}$ space because the variation in stellar radius within each spectral subtype is small enough that using $R_{p}/R_{\star}$ would introduce unnecessary redundancy. While both new planet candidates have juliet radius estimates nominally outside this range, their $1\sigma$ uncertainties are consistent with these limits. We factor this in by bootstrapping the radius and period uncertainties of the planets and candidates while estimating occurrence rates.
@@ 5.3 Completeness Grids
Figure 12: The completeness grid for our full sample of injected M-dwarfs, observed in both the TESS primary and first extended missions. Confirmed planets are indicated by blue circles, and the new planet candidate from this work is indicated by a white arrow.
@@ 5.4 Occurrence Rates
We estimated the occurrence rate for transiting GEMS within our sample for each of the three spectral subtype samples separately, using the Inverse Detection Efficiency Method (IDEM) described by Howard et al. (2012), and subsequently Christiansen et al. (2015) and others. Given the sparsity of planets, we did not estimate occurrences in each radius-period cell separately. Here, for the $n_{pl,j}$ planets in spectral subtype j, each planet $i$ contributes to the occurrence $f_{occ,j}$ as follows:
We accounted for errors in the period and radius estimates of all confirmed and candidate planets by bootstrapping the above equation across 5,000 iterations and accounting for the bootstrap-induced uncertainty ($\sigma_{\text{boot}}$). Our seven confirmed planets have a false positive probability (FPP) of 0, whereas for the new candidate lacking spectroscopic follow-up, we calculated occurrence rates under two extreme scenarios – assuming an FPP of 0 or 1. This approach encompasses all possible outcomes, providing a more comprehensive and transparent estimate than methods that rely on uncertain statistical validations alone. We note that bootstrap iterations in which a planet fell outside our injection bounds did not contribute to the occurrence rate. Since our new candidate has a median radius estimate outside these bounds, it did not significantly increase the occurrence rate even when we set $\text{FPP}_{\text{cand}}=0$. We present the occurrence rates calculated in this way as the “maximal” rates for completeness.
For the uncertainty in $f_{occ,j}$, we follow recommendations from Appendix A by Hsu et al. (2018), added in quadrature with errors introduced by bootstrapping:
Table 5: Occurrence rates of GEMS by spectral type
Type Median $\mathbf{M_{\star}~(M_{\odot})}$ $\mathbf{N_{\star}}$ $\mathbf{N_{p}}$ $\mathbf{N_{\rm cand}}$ Method $\mathbf{f_{\rm occ}}$    (nominal) [%] $\mathbf{f_{\rm occ}}$ (max) [%]
Early 0.54 27,995 2 0 IDEM $0.067\pm 0.047$  $\cdots$
Mid 0.35 37,948 4 1 IDEM $0.139\pm 0.069$ $0.139^{+0.072}_{-0.062}$
Late 0.18 79,669 1 0 IDEM $0.032\pm 0.032$  $\cdots$
Full 0.25 145,612 7 1 IDEM $065^{+0.025}_{-0.027} 0.0$$065^{+0.026}_{-0.026}$
Full 0.25 145,612 7 1 $\Gamma(1,1)^{*}$ $077\pm 0.027 0.0$$086\pm 0.029$
Full 0.25 145,612 7 1 $\Gamma(6.4,\,9763)^{*}$ $066\pm 0.018 0.0$$072\pm 0.019$
**footnotetext: Values shown only to illustrate the hierarchical Bayesian method; they are not adopted as final results in this work.
Note. — Each row lists the number of stars analysed ($N_{\star}$), confirmed GEMS ($N_{p}$), and new candidate GEMS ($N_{\rm cand}$). Occurrence rates are computed with the Inverse Detection Efficiency Method (IDEM) and, for comparison, a simplified Bayesian approach with a $\Gamma(\alpha_{0},\beta_{0})$ prior (see Section 5.4). The nominal rate counts only confirmed GEMS, while the “max” rate includes candidates and thus represents an upper limit. Stars lacking usable TGLC data are excluded from $N_{\star}$.
Figure 13: The occurrence rates of close-in giant planets around FGKM dwarfs based on RV searches (triangles), [^24] The inclusion of RV-based occurrence rates in this figure served purely as a qualitative comparison. We leave quantitative comparisons of RV- and transit-based occurrence rates to future work. transit searches (circles), and this work (filled crosses). “This Work (max)” refers to the maximal occurrence rate values including the new GEMS candidate (see Table 3); however, this is indistinguishable from the nominal value for the occurrence of GEMS around mid-type M-dwarfs. Upper limits are denoted by downward arrows and slight transparency, as in the case of Endl et al. (2006), Howard et al. (2012), and (Mignon et al., 2025). The upper limit from Endl et al. (2006) is shown at a representative stellar mass of 0.5 $M_{\odot}$, whereas those for Wright et al. (2012), Howard et al. (2012), and Wittenmyer et al. (2020) are shown at 1 $M_{\odot}$ given that the latter consist of a sample of FGK dwarfs. Finally, the dashed black line and associated grey error envelope represent an exponential fit to the black filled crosses from this work and the purple points from Beleznay & Kunimoto (2022). Fitting to these works allows us to span the stellar mass axis using only transit photometry based studies in TESS data within a similar period and radius range to our own work.
We note here that the contribution to the occurrence rate error from bootstrapping made up $1\%$ of the total error for early- and late-type bins, due to the small number of planets occurring on the edge of a given completeness bin (see Table 5). The occurrence rates for our three spectral classes, along with their median stellar masses, are summarized in Table 5. For mid M-dwarfs, we calculate a nominal occurrence rate with the confirmed planets, which increases to an upper bound when we include our newly identified GEMS candidate. The estimated upper bound assumes that this candidate is a genuine planet and hence brackets the range of current possibilities. Finally, we also combine all three spectral sub-types and find the occurrence rate of GEMS across our full sample of M-dwarfs, and list this in Table 5.
We note that while IDEM is computationally and intuitively simple, it may have limitations, especially in the regime of marginal detections where the transit S/N is comparable to the detection threshold, (i.e., for small planets, as has been highlighted by Hsu et al. (2018)). However, for large planets such as those under consideration here, the transit S/N is typically much larger than the detection threshold, as evidenced by our radius uncertainties in Table 3. Here, the biases induced by IDEM compared to other occurrence rate (and uncertainty) estimation techniques such as Approximate Bayesian Computing (ABC; see Figure 6 in Hsu et al., 2018) are expected to be smaller.

### v3

@@ 4.6 Stage 3: Observational and Literature Vetting
Table 2: The 23 objects which survived all auto and manual vetting prior to spectroscopic and literature validation. Of those 23, 7 are previously confirmed GEMS and 16 are astrophysical false positives (i.e., not transiting planets.) SB1/2/3 refer to single, double, or triple spectroscopic binaries.
@@
243921117 WASP 80 b Confirmed Early triaud_wasp_80b_2013
46432937 TIC 46432937 b Confirmed Early hartman_toi_762_2024
243641947 TOI-3235 b Confirmed Mid hobson_toi-3235_2023
335590096 TOI-4860 b Confirmed Mid almenara_toi-4860_2023
419411415 TOI-5205 b Confirmed Mid kanodia_toi-5205b_2023
178709444 TOI-762 A b Confirmed Mid hartman_toi_762_2024
67512645 TOI-6894 b Confirmed Late Bryant2025
@@
272519426 TOI-5693.01 BD Mid irwin_nltt_2010
381856446 WASP 145 A b (TIC 381856447 b) Background Planet Mid hellier_2019_wasp145ab
At this point, we referred back to our sample of known GEMS within 100 pc (see Section 2). We confirmed that we had successfully recovered all seven objects, reinforcing the robustness of our pipeline (see Table 2). Furthermore, we confirmed that we had recovered or reasonably discounted from our analysis all TOIs occurring within our input catalog and search bounds as of August 10, 2025. Only one TOI (5850, a spectroscopically confirmed single-lined binary), was removed prior to Stage 2, due to noisy observational data. This case, alongside other TOIs not explicitly mentioned in the main text of this work, is fully explained in Appendix Table 7.
We then consulted the TESS EB catalog (Prsa2022) as well as previous GEMS and EB surveys to determine the disposition of our remaining 16 candidates. These objects are listed in the bottom section of Table 2. We identified one candidate as an EB (Prsa2022) and one candidate as a BD (irwin_nltt_2010). We identified a further signal, ostensibly from TIC 381856446 / WASP 145 B, as a confirmed hot Jupiter orbiting the target’s much brighter K2 V companion, WASP 145 A (hellier_2019_wasp145ab). The remaining 13 candidates were subjected to spectroscopic validation using HPF, NEID and WINERED (see below for details on the use of these instruments.)
Spectroscopic vetting removed all new GEMS candidates, reducing our total GEMS sample from 20 to 7. This reduction of 65% highlights the importance of spectroscopic follow-up to eliminate FPs.
Table 3: Selected orbital parameters of all confirmed GEMS retrieved in our sample. Object parameters are quoted from the references in the final column. All stellar radii are taken from Mann15. Objects are sorted by $R_{p}$.
@@
178709444 TOI-762 A b $3.47\pm 7.2\cdot 10^{-7}$ $0.744\pm 0.017$ $0.7556^{+0.0057}_{-0.0067}$ $0.428$ hartman_toi_762_2024
335590096 TOI-4860 b $1.52\pm 3.0\cdot 10^{-7}$ $0.77\pm 0.03$ $0.29^{+0.06}_{-0.11}$ $0.355$ almenara_toi-4860_2023
243921117 WASP 80 b $3.07^{+8.3\cdot 10^{-7}}_{-7.9\cdot 10^{-7}}$ $0.999^{+0.030}_{-0.031}$ $0.215^{+0.020}_{-0.022}$ $0.636$ triaud_wasp_80b_2013
243641947 TOI-3235 b $2.59\pm 4.1\cdot 10^{-7}$ $1.02\pm 0.044$ $0.511^{+0.011}_{-0.012}$ $0.371$ hobson_toi-3235_2023
419411415 TOI-5205 b $1.63\pm 1.0\cdot 10^{-6}$ $1.03\pm 0.03$ $0.0016\pm 0.0002$ $0.400$ kanodia_toi-5205b_2023
46432937 TIC 46432937 b $44\pm 8.7\cdot 10^{-8} 1.1$$19\pm 0.030$ $0.825^{+0.0157}_{-0.0073}$ $0.539$ hartman_toi_762_2024
67512645 TOI-6894 b $3.37\pm 0.000296$ $1.33^{+0.57}_{-0.43}$ $0.927^{+0.551}_{-0.530}$ $0.220$ Bryant2025
@@ 4.6.1 HPF – Northern Hemisphere Targets I
For M-dwarfs in the northern hemisphere, we used the Habitable-zone Planet Finder (HPF; mahadevan_habitable-zone_2012; mahadevan_habitable-zone_2014) to obtain observations on our surviving candidates after the aforementioned checks. HPF is a near-infrared (8080127808080-12780 Å), environmentally stabilized (stefansson_versatile_16), fiber-fed (kanodia_overview_2018) spectrograph with laser-frequency comb calibration (Metcalf2019) on the 10 m Hobby-Eberly Telescope [^18] Based on observations obtained with the Hobby-Eberly Telescope (HET), which is a joint project of the University of Texas at Austin, the Pennsylvania State University, Ludwig-Ludwig-Maximilians-Universität München, and Georg-August Universität Gottingen. The HET is named in honor of its principal benefactors, William P. Hobby and Robert E. Eberly. (HET; ramsey_early_1998; hill_current_2012; hill_hetdex_2021) at the McDonald Observatory in West Texas, USA. We used the HxRGproc package (ninan_habitable-zone_2018) to correct for bias, non-linearity, cosmic rays and process the HPF slope images. We then used barycorrpy (kanodia_python_2018) to perform the barycentric correction on the individual spectra, which is the Python implementation of the algorithms from wright_barycentric_2014. Each visit consisted of two exposures of 969 seconds each that are subsequently combined by weighted averaging.
We followed the procedure described in Canas2023apogee to cross-correlate the observed spectra with synthetic spectra from the PHOENIX library (husser13) using the algorithms presented in Zucker2003. We only cross-correlate the HPF spectra in the echelle orders 464-6 ($\lambda\sim 8530-8890$ Å) and 141814-18 ($\lambda\sim 9930-10760$ Å) because of the minimal amount of telluric contamination in those regions. The resultant cross-correlation functions (CCFs) with HPF spectra are shown in Figure 7, and objects with multiple peaks are identified as FPs (hierarchical stellar systems). Using HPF spectra, we identified eight candidates that are spectroscopic binaries and are listed in Table 2. A detailed characterization of these false positives is beyond the scope of this manuscript and is left to future work (e.g., Boone2025).
Figure 7: CCF plots for the 13 TICs cited in Table 2 as astrophysical false positives by this work. Objects are arranged in the same order as Figure 6. The different colors represent distinct spectroscopic epochs, with CCF shifts indicating Doppler motion inconsistent with a planetary companion. In particular, the presence of multiple peaks or large velocity variations between observations reveal the target as an EB rather than a single-lined star hosting a planet.
@@ 4.6.2 NEID – Northern Hemisphere Targets II
We also used NEID, an environmentally stabilized (robertson_ultrastable_2019), fiber-fed (kanodia_stable_2023) high-resolution optical (3800 – 9300 Å) spectrograph (schwab_design_2016) on the 3.5 m WIYN telescope [^19] The WIYN Observatory is a joint facility of the NSF’s National Optical-Infrared Astronomy Research Laboratory, Indiana University, the University of Wisconsin-Madison, Pennsylvania State University, Purdue University and Princeton University. at Kitt Peak, Arizona, USA to spectroscopically validate some of our candidates. Each visit consisted of an 1800 s exposure using high-resolution mode ($R\sim$ 110,000). We used the Level 2 products from the NEID DRPv1.4.0202020https://neid.ipac.caltech.edu/docs/NEID-DRP/, which includes the wavelength-calibrated spectra. The NEID DRP also calculates the CCF for all objects by cross-correlating the observed spectra with binary masks (e.g., baranne_elodie_1996) that are customized to various spectral types (Bender2022). We used these CCFs to confirm that TIC-95112238 and TIC-60910638 were both spectroscopic double-lined binaries (SB2) with two peaks in the CCF for each star (see Figure 7).
@@ 5.1 Injection and Recovery
  • 0.7 $R_{J}$ $R_{\text{p, inj}}<1.5~R_{J}$, sampled in $R_{p}$ space. [^21] We chose to sample injections in $R_{p}$ as opposed to $R_{p}/R_{\star}$ space because the variation in stellar radius within each spectral subtype is small enough that using $R_{p}/R_{\star}$ would introduce unnecessary redundancy. Some GEMS fall near the edges of our injection bins, and we factor this in by bootstrapping across their radius and period uncertainties while estimating occurrence rates.
@@ 5.3 Completeness Grids
Figure 12: The completeness grid for our full sample of injected M-dwarfs, observed in both the TESS primary and first extended missions. Confirmed planets are indicated by blue circles.
@@ 5.4 Occurrence Rates
5.4.1 Bayesian Method